OpenStudy (akausen):

I need help with this problem. I have tried a few things, but really have no idea what to do. Thanks for any help. Consider the volume of the region shown below, which shows a right circular cone with top radius 3 cm and height 10 cm. We have used the notation Dy for Δy.

OpenStudy (anonymous):

post the pic & question

OpenStudy (compassionate):

@AravindG

OpenStudy (aravindg):

Can you tell me what will be f(x_k) here considering the fact that we are dealing with volumes?

OpenStudy (akausen):

I thought I would use the area of a circle, pi r^2, and multiply that by Dy, the width of each piece. That's all I've come up with.

OpenStudy (akausen):

I'm not sure what you mean by f(x_k) @AravindG

OpenStudy (aravindg):

Oh yes it's the area(my bad). f(x_k) is the area function of radius k.

OpenStudy (anonymous):

lol, sometimes math is so complicated and...WEIRD...

OpenStudy (akausen):

Oh, I see. I don't know where to go from there. I know the Riemann sum is the sum of f(xi)Dx where Dx= (b-a)/n= 10/n and xi=a+iDx=10i/n since a and b in the picture I provided are 0 and 10 respectively. I know those two are correct. However, I don't think I use this formula for this problem. I'm pretty sure I use pi r^2 somehow, but I'm not sure how to go about it. @AravindG